OpenStudy (anonymous):

write the following statement as a mathematical expression: "5 raised to the second power equals 25"

OpenStudy (nincompoop):

\[5^2=25\]

OpenStudy (nincompoop):

anything raised to the 2nd power means that the multiplicand and multiplier are the same 5 x 5 = 5^2 = 25 3 x 3 = 3^2 = 9 10 x 10 = 10^2 = 100 a x a = a^2

OpenStudy (nincompoop):

@idkatie please reply if you understand or have any question regarding the information I gave you.

OpenStudy (nincompoop):